• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Armstrong

    Log entry time 20:21:35 on July 9,2004

    Entry number 128055

    keyword=Shift plan

    
    

     HAPPEX-Hydrogen E99-115 Shift Plan: Friday, July 9 - Saturday July 10,  noon

    Updated: July 9, 20:15pm by   Dave Armstrong   Run Coordinator (pager 584-5411)


    Startup Program: 

    1. Check HRS magnet and septa current settings
    2. Check that High Voltages for HAPPEX detectors and lumi detectors are on.
    3. Check that the HRS detectors are off   (VDCs, etc.)
    4. As soon as CW beam returns:
      • Insert BeO target
      • Set Raster on
      • Ask for 5 µA on BeO target
      • Verify raster and beam location on BeO visually
      • Verify beam position with 0.3 mm in x,y of (0,0); if not, ask MCC to resteer.
      • With spot++, verify the raster size. It should be 3.5x3.5 mm2

    Branch Point:

    If at this stage it is before midnight or after 4 AM, page Kent and Lisa, and have them come in to set up the helicity-control systems: Hall A and Hall C IA cells, PITA voltages, RHWP.

    If it is between midnight and 4 AM, proceed as below:

    1. Move hydrogen target in to beam (loop 2, 20 cm racetrack cell)
    2. Ensure 60 Hz fan speed.
    3. Make sure Fast Feedback is on (MCC)
    4. Ask for 20 µA on target, with raster on.
    5. Check that target is OK
    6. Start HAPPEX DAQ run (not yet production running)
    7. Watch Septa temperatures. Record equilibrium beam-on values (may take 20 minutes to settle)
    8. Meanwhile:
      • Look through Panguin plots - verify detectors, lumis, BPMs, BCMs... Lumi's and detectors all should be at around 30 K channels.
      • turn on beam modulation (dithering). Verify that typical plot produced.
      • take profile scanner run; verify that the profiles are close to the standard.
    9. Calculate "safe" running beam current (such that max. septum coil temps are less than 7.75 K)
    10. end run, ask for beam current at the maximum safe current.
    11. Start new HAPPEX DAQ run
    12. Verify septa temps. are stabilized; don't allow them to exceed 7.75 K! Turn beam current down if needed. Record septa. stable temps and beam current.
    13. Try to make Compton measurement. Expected value with this gun is around 76%.

    Production settings

    Tasks

    General Tasks:

    Please log below (and in the HALOG) any unusual observations and problems solved
    Date/Time Name Comment HALOG